A cockroach of mass M/2 starts moving with velocity v on the circumference of a disc of mass M and R what will be the angular velocity of the disc?

Grade:12th pass

1 Answers

Umme
5 years ago
First draw the free body diagram of the cockroach 
N=Mg/2
f=Mn^2R/2
Where n =angular velocity
f=friction 
n=√ng/R